In a shopping mall an average of 3 out of every 5 customers use visa credit...

  1. In a shopping mall an average of 3 out of every 5 customers use visa credit cards. A sample of 10 customers is selected.
    1. Find the probability that exactly 6 customers use visa credit card.
    2. Find the probability that less than 9 customers use visa credit card.
    3. Now assume that a sample of 50 customers are selected. Find the smallest value of n such that P(X<n) ≥ 0.9, where the random variable X represents the number of customers who uses visa credit card.
Answer #1

let xbe randon variable denotes The number of customers who use visa credit cards. probability of having credit card, p= 3 = 0.6 Somple size, n=10 X. Binomial (n=10, p=0.6) The probability man function of xis, P(x= x) = (0) (0.65 Ciro.g; x x=0,1,2... 10 as The probability. That exactly 6 customers use credit card: (X=6) = (10) 10.63 (-0.60- - 0.250822656 0.2508 The probability that lem Than a customers use visa credit card : P(X 29. - [P(x39]] = [p(x=9) + P(x=10)] [() 60-03.10.+C)(0.2) cons]